# FOGNA: An effective Sum-Difference Co-Array Design Based on Fourth-Order Cumulants

### v2

@@ I Introduction
Minimum redundancy array (MRA) [19] is a foundational structure in order to obtain as large DOF as possible by reducing redundant sensors. However, the non-closed form expressions for sensor positions hinder MRA’s scalability and complicate large-scale array design. This limitation leads to extensive research on nested arrays [13] and coprime arrays [14], which offer significant advantages with their closed-form expressions for the sensor locations. The success of nested arrays and coprime arrays inspires further developments aimed at enhancing DOF, including augmented coprime arrays [20], enhanced nested arrays [21] and arrays based on the maximum element spacing criterion [22]. With the respect of DOA estimation, traditional subspace-based methods [24] only utilize the consecutive lags of the DCA, making a hole-free configuration advantageous. Consequently, hole-filling strategies have been proposed to create new coprime arrays-like with a hole-free DCA [25], [26]. Obviously, various DCAs based on SOC have been widely studied in DOA estimation because of its significantly enhanced DOF [13], [14], [21], [22], [23].
@@ III-B The FOECA With Mutual Coupling
$\boldsymbol{\tilde{z}}_{\boldsymbol{x}}^{(1)}=\boldsymbol{C}^{(1)}_{vec}\boldsymbol{B}^{(1)}\boldsymbol{p}_{\boldsymbol{s}}^{(1)},$ (28a)
$\boldsymbol{\tilde{z}}_{\boldsymbol{x}}^{(2)}=\boldsymbol{C}^{(2)}_{vec}\boldsymbol{B}^{(2)}\boldsymbol{p}_{\boldsymbol{s}}^{(2)},$ (28b)
$\boldsymbol{\tilde{z}}_{\boldsymbol{x}}^{(3)}=\boldsymbol{C}^{(3)}_{vec}\boldsymbol{B}^{(3)}\boldsymbol{p}_{\boldsymbol{s}}^{(3)}.$ (28c)
@@
$\boldsymbol{\tilde{z}}_{\boldsymbol{x}}$ $=[{\boldsymbol{\tilde{z}}_{\boldsymbol{x}}^{(1)}}^{T},{\boldsymbol{\tilde{z}}_{\boldsymbol{x}}^{(2)}}^{T},{\boldsymbol{\tilde{z}}_{\boldsymbol{x}}^{(3)}}^{T}]^{T}$ (29)
$\triangleq\boldsymbol{\tilde{C}}_{vec}\boldsymbol{B}\boldsymbol{p}_{s}\in\mathbb{C}^{3N^{4}\times 1},$
@@ IV-C The Maximum DOF of FOGNA With the Given Number of Physical Sensors
[…]
$=2(2(\lfloor\frac{2(N-N_{1})+1}{4}\rfloor)+1)(2E_{1}$
@@
    Array config. DOF
    FL-NA $\mathcal{O}(\frac{N^{4}}{128})$
    SE-FL-NA $\mathcal{O}(\frac{N^{4}}{64})$
    FO-Fractal $\mathcal{O}(\frac{N^{4}}{32})$
    FOGNA $\mathcal{O}(\frac{N^{4}}{2})$
@@ V-A Comparison of the DOF for Different Arrays
We compare the DOF of the proposed method with those of FL-NA [28], SE-FL-NA [30], FOFractal (NA) [44] and SD-FODC (NA) [18] for given the fixed number of physical sensors, where FL-NA and SE-FL-NA adopt the array structure for obtaining maximum DOF and the nested array is used as the basic array of FO-Fractal and SD-FODC. The comparing results are listed in Table II and the variations of DOF versus the number of sensors for five methods are shown in Fig. 3. The results in Table II show that the DOF of FOGNA have significantly increased compared to other four methods under the different given number of physical sensors. In addition, it can be seen that as the number of sensors $N$ increases, the DOF growth rate of the proposed FOGNA is faster than those of the other four FODCAs in Fig. 3, and the advantages of the proposed array are more obvious with a large number of physical sensors.
@@ V-B Mutual Coupling Matrices
Array config.    FL-NA   SE-FL-NA   FO-Fractal (NA)   SD-FODC (NA)   FOGNA
@@ Appendix A Proof of the Corollary 1
$(2E_{1}+1)\}+1,$
$E_{1}\leq$ $(N_{1}-1)\lceil\frac{N_{1}-1}{4}\rfloor+N_{1}-1$
$\leq$ $(N-1)\lceil\frac{N-1}{4}\rfloor+N-1,$

### v3

@@ I Introduction
Minimum redundancy array (MRA) [18] is a foundational structure in order to obtain as large DOF as possible by reducing redundant sensors. However, the non-closed form expressions for sensor positions hinder MRA’s scalability and complicate large-scale array design. This limitation leads to extensive research on nested arrays [13] and coprime arrays [14], which offer significant advantages with their closed-form expressions for the sensor locations. The success of nested arrays and coprime arrays inspires further developments aimed at enhancing DOF, including augmented coprime arrays [19], enhanced nested arrays [20] and arrays based on the maximum element spacing criterion [21]. With the respect of DOA estimation, traditional subspace-based methods [23] only utilize the consecutive lags of the DCA, making a hole-free configuration advantageous. Consequently, hole-filling strategies have been proposed to create new coprime arrays-like with a hole-free DCA [24].Obviously, various DCAs based on SOC have been widely studied in DOA estimation because of its significantly enhanced DOF [13], [14], [20], [21], [22].
@@ III-B The FOECA With Mutual Coupling
$\boldsymbol{\tilde{z}}_{\boldsymbol{x}}^{(1)}=\boldsymbol{C}^{(1)}_{vec}\boldsymbol{B}^{(1)}\boldsymbol{p}_{\boldsymbol{s}}^{(1)},$ (28)
$\boldsymbol{\tilde{z}}_{\boldsymbol{x}}^{(2)}=\boldsymbol{C}^{(2)}_{vec}\boldsymbol{B}^{(2)}\boldsymbol{p}_{\boldsymbol{s}}^{(2)},$
$\boldsymbol{\tilde{z}}_{\boldsymbol{x}}^{(3)}=\boldsymbol{C}^{(3)}_{vec}\boldsymbol{B}^{(3)}\boldsymbol{p}_{\boldsymbol{s}}^{(3)}.$
@@
$\boldsymbol{\tilde{z}}_{\boldsymbol{x}}$ $=[{\boldsymbol{\tilde{z}}_{\boldsymbol{x}}^{(1)}}^{T},{\boldsymbol{\tilde{z}}_{\boldsymbol{x}}^{(2)}}^{T},{\boldsymbol{\tilde{z}}_{\boldsymbol{x}}^{(3)}}^{T}]^{T}\triangleq\boldsymbol{\tilde{C}}_{vec}\boldsymbol{B}\boldsymbol{p}_{s}\in\mathbb{C}^{3N^{4}\times 1},$ (29)
@@ IV-C The Maximum DOF of FOGNA With the Given Number of Physical Sensors
$h(N_{1})$ $\triangleq\text{DOF}=2(2(\lfloor\frac{2(N-N_{1})+1}{4}\rfloor)+1)(2E_{1}$ (44)
@@
Array config. FL-NA SE-FL-NA FO-Fractal FOGNA
DOF $\mathcal{O}(\frac{N^{4}}{128})$ $\mathcal{O}(\frac{N^{4}}{64})$ $\mathcal{O}(\frac{N^{4}}{32})$ $\mathcal{O}(\frac{N^{4}}{2})$
@@ V-A Comparison of the DOF for Different Arrays
We compare the DOF of the proposed method with those of FL-NA [26], SE-FL-NA [28], FO-Fractal (NA) [40] and SD-FODC (NA) [17] for given the fixed number of physical sensors, where FL-NA and SE-FL-NA adopt the array structure for obtaining maximum DOF and the nested array is used as the basic array of FO-Fractal and SD-FODC. The comparing results are listed in Table II and the variations of DOF versus the number of sensors for five methods are shown in Fig. 3. The results in Table II show that the DOF of FOGNA have significantly increased compared to other four methods under the different given number of physical sensors. In addition, it can be seen that as the number of sensors $N$ increases, the DOF growth rate of the proposed FOGNA is faster than those of the other four FODCAs in Fig. 3, and the advantages of the proposed array are more obvious with a large number of physical sensors.
@@ V-B Mutual Coupling Matrices
Array FL-NA SE-FL FO-Fractal SD-FODC FOGNA
config. -NA (NA) (NA)
@@ Appendix A Proof of the Corollary 1
$E_{1}\leq$ $(N-1)\lceil\frac{N-1}{4}\rfloor+N-1,$